# DUET: Distilled LLM Unlearning from an Efficiently Contextualized Teacher

### v1

@@ 4.4 Performance Overview
Harry Potter (MUSE-Books). Table 1 reports overall results for the Llama 3.2-3B-Instruct LLM on the Harry Potter (HP) benchmark. DUET demonstrates competitive or higher performance compared with state-of-the-art unlearning methods, which effectively removes undesirable knowledge while preserving general knowledge utility. Specifically, GA unlearns the knowledge of HP at the consequences of catastrophic forgetting. On the other hand, augmenting a retention loss to GA can mitigate utility drop, yet hurt the unlearning effectiveness. Similar phenomena were observed on NPO. In contrast, DUET maintains the highest general utility preservation, while achieving more effective unlearning than methods such as NPO or its variants. Most baselines are sensitive to the size and format of forgetting data, whereas our method can benefit from a lightweight dataset ${\mathcal{D}}_{f}^{\text{query}}$ (Sec 4.1), owing to its fine-grained knowledge distillation design.
@@
WMDP (Cyber/Bio). As shown in Table 2, GA and FLAT shows a catastrophic utility drop, while methods such as Refusal Training or GA combined with a KL regularization showed marginal effects on the forgetting domain. While most methods struggle to balance unlearning and retention and often sacrifice one for the other, our method notably delivers the best overall performance shifts, followed by RMU as the closest competitor, a method carefully tailored for WMDP benchmarks.
@@ 7 Reproducibility Statement
We aim to make all results fully reproducible. An anonymized repository will be released at https://anonymous.4open.science/r/DUET-FAF3 containing all source code and the complete datasets (training/retain/forget splits and test sets), along with scripts to regenerate every table and figure end-to-end. Dataset construction and evaluation splits are summarized in Sec 4.1; baseline implementations and protocols are in Sec 4.2 and 4.3; the training procedure is outlined in Alg 1. The appendix provides the full hyperparameter and training specifications in A.3, including the exact teacher prefixes A.1, training temperature, number of epochs, batch size, learning rate, environment details, so that results can be reproduced precisely. As shown in Table A.4, our method exhibits a clear trade-off between forgetting (R-Forget500) and retention (R-Retain) across different learning rates, epochs, and schedulers.
@@ 8 LLM Usage Statement
Beyond serving as the subjects of study in our experiments (i.e., LLMs we trained and evaluated, and used to draft candidate teacher prefixes), we also used an off-the-shelf general-purpose LLM (GPT-5) as a writing assistant. Specifically, it was employed to (i) polish grammar and wording and (ii) suggest LaTeX formatting adjustments for figures and tables (e.g., column widths, wrap/wrapfig usage, captions). The LLM did not generate research ideas, models, analyses, or claims; all technical design, data curation, and conclusions are by the authors. All suggested text/formatting was reviewed and edited by the authors, and no non-public or sensitive data were shared with the LLM.
@@ A.4 Additional Results on HP QA (Hyperparameter Sweep)
We evaluate different training hyperparameters on the HP (Harry Potter) QA subset.
@@ A.5 Additional Results on HP (More Baselines)
To complement the main results, we additionally include SFR Huang et al. (2024) and UNDIAL Dong et al. (2024) as baselines for comparative analysis. SFR, which operates by following a Hessian-guided forgetting direction on a remain-preserving manifold, and UNDIAL, a distillation-based approach that directly shifts logits toward predefined refusal tokens, represent two conceptually distinct families of unlearning techniques. The complete experimental results after adding these two baselines are shown in the Table  10.
@@
As reflected in Table  10, SFR does not outperform our approach and exhibits significant forgetting of general knowledge, indicating that its manifold-guided update direction does not effectively decouple targeted forgetting from utility preservation.
@@ A.6 Distillation from Gradient Ascent Models
A.7 Distillation from Multiple Unlearning Teachers
(1) Distillation from multiple heterogeneous teachers.
(2) Distillation from multiple prefix-guided teachers.
  1. 1.
  2. 2.
  3. 3.
base model 32.13 39.99 84.29 61.46 0.00
DUET 4.27 4.86 78.33 61.52 57.09
Plan A (GA, NPO, one prefix) 7.94 10.28 58.58 45.43 12.16
Plan A (GA, NPO, one prefix; retain free) 13.98 14.92 82.83 61.04 41.34
Plan B (3 different prefixes) 8.32 9.77 81.02 60.13 49.43
A.8 Generalization Across LLMs
@@
A.9 Ablation on Multi-Step Decoding Distillation
@@
A.10 Ablation on Training Data Size
@@ Results.
A.11 Adversarial Evaluation with Jailbreak Attacks
@@
A.12 Ablation on Prefix Usage in Retention Data

### v2

@@ 4.4 Performance Overview
Harry Potter (MUSE-Books). Table 1 reports overall results for the Llama 3.2-3B-Instruct LLM on the Harry Potter (HP) benchmark. DUET demonstrates competitive or higher performance compared with state-of-the-art unlearning methods, which effectively removes undesirable knowledge while preserving general knowledge utility. Specifically, GA unlearns the knowledge of HP at the cost of catastrophic forgetting. On the other hand, augmenting a retention loss to GA can mitigate utility drop, yet hurt the unlearning effectiveness. Similar phenomena were observed on NPO. In contrast, DUET maintains the highest general utility preservation, while achieving more effective unlearning than methods such as NPO or its variants. Most baselines are sensitive to the size and format of forgetting data, whereas our method can benefit from a lightweight dataset ${\mathcal{D}}_{f}^{\text{query}}$ (Sec 4.1), owing to its fine-grained knowledge distillation design.
@@
WMDP (Cyber/Bio). As shown in Table 2, GA and FLAT show a catastrophic utility drop, while methods such as Refusal Training or GA combined with a KL regularization showed marginal effects on the forgetting domain. While most methods struggle to balance unlearning and retention and often sacrifice one for the other, our method notably delivers the best overall performance shifts, followed by RMU as the closest competitor, a method carefully tailored for WMDP benchmarks.
@@ 4.6.3 Robustness of DUET on the number of Top Candidate Logits
4.6.4 Distillation from Multiple Unlearning Teachers
(1) Distillation from multiple heterogeneous teachers.
(2) Distillation from multiple prefix-guided teachers.
  1. 1.
  2. 2.
  3. 3.
base model 32.13 39.99 84.29 61.46 0.00
DUET 4.27 4.86 78.33 61.52 57.09
Plan A (GA, NPO, one prefix) 7.94 10.28 58.58 45.43 12.16
Plan A (GA, NPO, one prefix; retain free) 13.98 14.92 82.83 61.04 41.34
Plan B (3 different prefixes) 8.32 9.77 81.02 60.13 49.43
@@ 5 Conclusion
Acknowledgments
This work was supported in part by the NAIRR Pilot (NAIRR250140) and an NVIDIA Academic Grant.
@@ 7 Reproducibility Statement
We aim to make all results fully reproducible. An anonymized repository will be released at https://github.com/EasonZhong99/DUET containing all source code and the complete datasets (training/retain/forget splits and test sets), along with scripts to regenerate every table and figure end-to-end. Dataset construction and evaluation splits are summarized in Sec 4.1; baseline implementations and protocols are in Sec 4.2 and 4.3; the training procedure is outlined in Alg 1. The appendix provides the full hyperparameter and training specifications in A.3, including the exact teacher prefixes A.1, training temperature, number of epochs, batch size, learning rate, environment details, so that results can be reproduced precisely. As shown in Table A.4, our method exhibits a clear trade-off between forgetting (R-Forget500) and retention (R-Retain) across different learning rates, epochs, and schedulers.
@@ 8 LLM Usage Statement
Beyond serving as the subjects of study in our experiments (i.e., LLMs we trained and evaluated, and used to draft candidate teacher prefixes), we also used an off-the-shelf general-purpose LLM (GPT-5) as a writing assistant. Specifically, it was employed to (i) polish grammar and wording and (ii) suggest LaTeX formatting adjustments for figures and tables (e.g., column widths, wrap/wrapfig usage, captions). The LLM did not generate research ideas, models, analyses, or claims; all technical design, data curation, and conclusions are by the authors. All suggested text/formatting were reviewed and edited by the authors, and no non-public or sensitive data were shared with the LLM.
@@ A.4 Additional Results on HP QA (Hyperparameter Sweep)
We evaluate different training hyperparameters on the HP (Harry Potter) QA subset, as shown in Table 10.
@@ A.5 Additional Results on HP (More Baselines)
To complement the main results, we additionally include SFR Huang et al. (2024) and UNDIAL Dong et al. (2025) as baselines for comparative analysis. SFR, which operates by following a Hessian-guided forgetting direction on a remain-preserving manifold, and UNDIAL, a distillation-based approach that directly shifts logits toward predefined refusal tokens, represent two conceptually distinct families of unlearning techniques. The complete experimental results after adding these two baselines are shown in Table 11.
Overall, Table 11 indicates that DUET achieves the most balanced performance by jointly improving forgetting while preserving utility.
As reflected in Table  11, SFR does not outperform our approach and exhibits significant forgetting of general knowledge, indicating that its manifold-guided update direction does not effectively decouple targeted forgetting from utility preservation.
@@ A.6 Distillation from Gradient Ascent Models
A.7 Generalization Across LLMs
@@
A.8 Ablation on Multi-Step Decoding Distillation
@@
A.9 Ablation on Training Data Size
@@ Results.
A.10 Adversarial Evaluation with Jailbreak Attacks
@@
A.11 Ablation on Prefix Usage in Retention Data